OpenStudy (anonymous):

if i solve for x (3x+8) = (5x-20) will the answer be 20?

jimthompson5910 (jim_thompson5910):

(3x+8) = (5x-20) 3x+8 = 5x-20 3x+8 - 5x = -20 3x - 5x = -20-8 -2x = -28 x = -28/(-2) x = 14

OpenStudy (anonymous):

thats wrong..

jimthompson5910 (jim_thompson5910):

hmm are there fractions outside of the parenthesis?

jimthompson5910 (jim_thompson5910):

seems odd how there are parenthesis when they're not really needed so that's why I'm thinking something is missing

OpenStudy (luigi0210):

Are you sure? @smarie96

OpenStudy (anonymous):

yeah its 14 :)

OpenStudy (anonymous):

well my online robot teacher said it was 50

jimthompson5910 (jim_thompson5910):

does not compute...hmm can you do a screenshot?

jimthompson5910 (jim_thompson5910):

so you can show us the full problem

OpenStudy (anonymous):

3x + 8 = 5x - 20 3x - 5x = -20 - 8 -2x = - 28 x = 14

OpenStudy (luigi0210):

o.O

OpenStudy (anonymous):

it's a robot.. robot do mistakes :) but look :) http://www.wolframalpha.com/input/?i=%283x+%2B+8%29+%3D+%285x++-+20%29+solve+for+x

OpenStudy (anonymous):

wait a minute.....she has asked a lot of these questions.....she needs to learn how to do them so she doesn't have to ask the same questions

OpenStudy (anonymous):

because right now she is on another questionthat she asked...the same type